The mean water consumption of 10 students in a week is 28 liters with a standard deviation of 1.6 liters. Consider 95% confidence level, the upper limit is


1
Expert's answer
2022-04-27T00:29:49-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 9" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.262156."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(28-2.262156\times\dfrac{1.6}{\sqrt{10}}, 28+2.262156\times\dfrac{1.6}{\sqrt{10}})"

"=(26.85543, 29.14457)"

The upper limit is 29.14457 liters.